OpenStudy (anonymous):

FInd an equation of a line that goes through (5/8,2/9) and is vertical.

OpenStudy (anonymous):

SO then 5/8 has to stay the same?

OpenStudy (anonymous):

Sorry my mistake it's not going to be a function where you can plot. A function only has one distinct value for each x.

OpenStudy (anonymous):

So that can be a point but you can't find a function.

OpenStudy (anonymous):

What?

OpenStudy (anonymous):

You can't have a vertical function.

OpenStudy (anonymous):

I get that... But you have confused me with everything else.

OpenStudy (anonymous):

x=5/8

OpenStudy (anonymous):

vertical line x = number in your case it is \(x=\frac{5}{8}\)

OpenStudy (anonymous):

How would you set up the equation to get the answer.. That is what I am confused on.

OpenStudy (lgbasallote):

maybe drawing can help... so...x = 5/8

OpenStudy (anonymous):

vertical line means x is the same , no matter what the y is

OpenStudy (lgbasallote):

the equation is x = 5/8 itself

OpenStudy (anonymous):

But how? I guess I just dont understand what you did with the 2/9.. You set it to 0?

OpenStudy (anonymous):

Yes x=5/8 is the equation but you can't come up with a function where you can plug in any value of X and come up with a different value of Y

OpenStudy (anonymous):

So if you cant come up with a function you would have x=5/8 but to get that you would 0(2/9)?